# Identifying and Solving Conditional Image Leakage in Image-to-Video Diffusion Model

### v1

@@ Abstract
Diffusion models have obtained substantial progress in image-to-video (I2V) generation. However, such models are not fully understood. In this paper, we report a significant but previously overlooked issue in I2V diffusion models (I2V-DMs), namely, conditional image leakage. I2V-DMs tend to over-rely on the conditional image at large time steps, neglecting the crucial task of predicting the clean video from noisy inputs, which results in videos lacking dynamic and vivid motion. We further address this challenge from both inference and training aspects by presenting plug-and-play strategies accordingly. First, we introduce a training-free inference strategy that starts the generation process from an earlier time step to avoid the unreliable late-time steps of I2V-DMs, as well as an initial noise distribution with optimal analytic expressions (Analytic-Init) by minimizing the KL divergence between it and the actual marginal distribution to effectively bridge the training-inference gap. Second, to mitigate conditional image leakage during training, we design a time-dependent noise distribution for the conditional image, which favors high noise levels at large time steps to sufficiently interfere with the conditional image. We validate these strategies on various I2V-DMs using our collected open-domain image benchmark and the UCF101 dataset. Extensive results demonstrate that our methods outperform baselines by producing videos with more dynamic and natural motion without compromising image alignment and temporal consistency. The project page: https://cond-image-leak.github.io/.
@@ 1 Introduction
Image-to-video (I2V) generation aims to generate videos with dynamic and natural motion while maintaining the appearance of the given image. I2V generation allows users to guide video creation from the input image (and optional text), thus increasing controllability and flexibility in content creation. Like the remarkable progress in text-to-image (T2I) generation  [40, 38, 20, 8, 18, 5] and text-to-video (T2V) generation  [10, 42, 6, 11], diffusion models have also obtained promising results for I2V generation [9, 15, 54, 12, 58]. However, such models are not fully understood.
Figure 1: Schematic diagram of conditional image leakage. Ideally, I2V-DMs predict clean videos from noisy inputs, using the conditional image as auxiliary content guidance. However, at large time steps, the heavily corrupted input retains minimal video detail, causing the model tends to over-rely on the detailed conditional image and neglect the crucial task of synthesizing video from noisy inputs.
@@
In this paper, we identify a significant but previously overlooked issue in the image-to-video diffusion models (I2V-DMs), namely, conditional image leakage (see Sec. 3.1). Ideally, the I2V-DMs should predict the added noise or clean video mainly from noisy inputs, with the conditional image serving as auxiliary guidance. However, as the diffusion process progresses—especially at large time steps, the heavily corrupted input retains minimal video detail, while the conditional image preserves extensive detail of the target video. This biases the model to over-rely on the conditional image while neglecting the essential task of synthesizing video from noisy inputs, leading to videos lacking dynamic and vivid motion. For instance, I2V-DMs that take a shortcut by directly replicating the conditional image may even achieve a lower loss. To validate this issue, we corrupt a ground truth (GT) clean video data using the forward transition kernel and compute one-step clean video predictions across all time steps. As illustrated in Fig. 2, predicted clean videos exhibit markedly reduced motion than GT at large time steps, with videos becoming almost entirely static at the final step.
@@
Based on the analysis, we attempt to address the challenge in both inference and training. First, we present a simple yet effective training-free inference strategy, where we start the video generation process from an earlier time step instead of the final one, thus avoiding the unreliable late-time steps of the I2V-DMs (see Sec. 4.1). To further enhance performance, we derive an initial noise distribution with optimal analytic expressions (Analytic-Init) by minimizing the KL divergence between it and the actual marginal distribution to bridge the training-inference gap. Second, to mitigate the conditional image leakage during training, we introduce a time-dependent noise distribution (TimeNoise) that interferes with the conditional image, particularly at large time steps (Sec. 4.2). In principle, the distribution should favor high noise levels at large time steps to sufficiently disrupt the conditional image, then gradually decrease the noise levels as the time step diminishes. To achieve this, we utilize a logit-normal distribution with a center that monotonically increases over time and conduct a systematic analysis of its two hyperparameters. Notably, our strategies are designed to be plug-and-play, making them adaptable to various I2V-DMs based on both VP-SDE [12, 54] and VE-SDE [9] frameworks. Finally, the clean videos predicted by our method maintain motion dynamics comparable to the GT across all time steps, indicating the absence of conditional image leakage (see Fig. 6).
@@
Empirically, we validate our methods on various I2V-DMs [9, 54, 12] using our collected open-domain images (ImageBench) and the UCF101 dataset. We conduct a user study on the ImageBench and report FVD [49] and IS [7] on UCF101. Extensive experimental results demonstrate that our strategies outperform baselines by producing videos with markedly more dynamic and natural motion, all while maintaining image alignment and temporal consistency.
@@ 2 Background
Diffusion models gradually perturb data ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$ through a forward diffusion process and reserve it to recover the data. The forward transitional kernel $q_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})$ is given by
@@
where $\alpha_{t}$ and $\sigma_{t}$ are the noise schedule and chosen to ensure that ${\bm{x}}_{T}$ contains minimal information about ${\bm{x}}_{0}$. Two prevalent types of forward stochastic differential equation (SDEs) are commonly used [44, 28]. One is the variance-preserving SDE (VP-SDE) [40, 23], where $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ with $\alpha_{t}\rightarrow 0$ as $t\rightarrow T$, ensuring $p_{T}({\bm{x}}_{T})=\mathcal{N}({\bm{0}},{\bm{I}})$. The other is the variance exploding SDE (VE-SDE), where $\alpha_{t}=1$ and $\sigma_{T}$ is set to a large constant, resulting in $p_{T}({\bm{x}}_{T})={\mathcal{N}}({\bm{0}},\sigma_{T}^{2}{\bm{I}})$. Such models can be parameterized using a noise-prediction model $\epsilon_{\theta}({\bm{x}}_{t},t)$ (${\bm{\epsilon}}$-prediction) [23] to optimize:
@@
where the noisy input ${\bm{x}}_{t}\sim q_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})$. Alternative parametrizations such as ${\bm{x}}_{0}$-prediction [28], ${\bm{v}}$-prediction [41], $F$-prediction [27] are also commonly used. The ${\bm{\epsilon}}$-prediction and ${\bm{x}}_{0}$-prediction loss aims to predict the added noise or clean video from noisy input ${\bm{x}}_{t}$. Starting from ${\bm{x}}_{T}\sim p_{T}({\bm{x}}_{T})$, various samplers [43, 32, 33, 27] can be employed to generate data.
@@
Given any image ${\bm{y}}_{0}$ from the open domain, the goal of I2V is to generate a video $X_{0}=\{{\bm{x}}_{0}^{i}\}_{i=1}^{N}$ with dynamic and natural motion while maintaining the appearance of ${\bm{y}}_{0}$. This task can be formulated as designing a conditional distribution $p_{\theta}(X_{0}|{\bm{y}}_{0})$, which be achieved by a conditional diffusion model:
@@
where $X_{t}\sim q_{t|0}(X_{t}|X_{0})$. Typically, ${\bm{y}}_{0}$ is the first frame of $X_{0}$ and DynamiCrafter [54] adopts a randomly selected frame from $X_{0}$ as ${\bm{y}}_{0}$. The key issue is effectively integrating the conditional image ${\bm{y}}_{0}$ into the diffusion model. Most methods use CLIP image embeddings [39] to maintain the semantic content of ${\bm{y}}_{0}$. Notably, VideoCrafter1 [12] and Dynamicrafter [54] employ the last layer’s full patch visual tokens from the CLIP ViT, enriching the encoded information, and other approaches prefer the class token layer. Yet, solely depending on these embeddings, such as in VideoCrafter1 [12], compromises detail retention, impacting image alignment. To enhance detail representation, I2VGen-XL [58] combines the conditional image with the noisy initial frame, while VideoComposer [51] develops a STC-encoder for multiple conditions. Although superior to CLIP image embeddings, these strategies still fail to fully retain the conditional image content. To mitigate this, AnimateAnything [15], Dynamicrafter [54] and SVD [9] choose to directly concatenate noisy video $X_{t}$ with ${\bm{y}}_{0}$, which provides detailed information.
Figure 2: Identifying conditional image leakage in I2V-DMs. We corrupt a ground truth (GT) clean video and use it to compute $\hat{X}_{t\rightarrow 0}$ prediction across time steps. As time progresses, $\hat{X}_{t\rightarrow 0}$ exhibits markedly reduced motion than GT, with the videos becoming nearly static at the final step $T$, suggesting the existence of conditional image leakage in a large time step.
3 Conditional Image Leakage
In this section, we first identify a commonly overlooked issue in I2V-DMs: conditional image leakage, discussed in Sec. 3.1. We then discuss existing I2V-DMs with respect to this issue in Sec. 3.2.
@@ 3.1 Identifying Conditional Image Leakage in Image-to-video Diffusion Models
Although existing I2V-DMs discussed in Sec. 2 have obtained significant progress, such models are not fully understood. This section revisits and underscores a critical yet previously overlooked issue: conditional image leakage (CIL). As shown in Fig. 1, ideally, I2V-DMs should predict the added noise ($\epsilon$-prediction) or clean video (${\bm{x}}_{0}$-prediction) mainly from noisy input $X_{t}$, using the conditional image ${\bm{y}}_{0}$ as auxiliary content guidance. However, as the diffusion process progresses—particularly at large time steps, the heavily corrupted input $X_{t}$ contains minimal information about $X_{0}$, making it challenging to predict clean frames. Meanwhile, the conditional image retains considerable detail of $X_{0}$. This biases the model to over-rely on the conditional image while neglecting the essential task of synthesizing video from noisy inputs. For instance, a diffusion model that takes a shortcut by directly replicating ${\bm{y}}_{0}$ can even achieve a lower loss than one that accurately predicts the clean video from a heavily corrupted $X_{t}$. As a result, the synthesized videos lack dynamic and vivid motion. Notably, recent strategies that involve adjusting the noise schedule towards more noise [14, 26, 9, 31] may further exacerbate this issue (see details in Appendix C).
To validate this issue, we corrupt a ground truth (GT) clean video $X_{0}$ using the forward transition kernel in Eq. (1) and use it as the noisy input to compute the one-step prediction $\hat{X}_{t\rightarrow 0}$ at time $t$:
@@
Ideally, $\hat{X}_{t\rightarrow 0}$ attempts to predict the GT $X_{0}$ from noisy input $X_{t}$ and exhibit comparable motion dynamics. However, as shown in Fig. 2, as time progresses—particularly at large time steps, $\hat{X}_{t\rightarrow 0}$ exhibits markedly reduced motion than GT, with the videos becoming nearly static at the final step $T$, suggesting the existence of the conditional image leakage. Consequently, it will generate an almost still video starting from $T$ by duplicating the conditional image (see Fig. 6).
Figure 3: Understanding exiting work from conditional image leakage. (a) I2VGen-XL [58] mitigates the leakage at the expense of image alignment. (b) Increasing the SVD motion score induces unnatural camera movements with stationary objects. Our inference strategy, however, produces dynamic and natural motion in both people and the teapot.
3.2 Understanding Existing Work from Conditional Image Leakage
We now examine three popular I2V-DMs with respect to conditional image leakage: I2VGEN-XL [58], Dynamicrafter [54], and SVD [9]. Although these models do not explicitly address conditional image leakage, we believe their strategies may mitigate it to some extent.
Firstly, I2VGEN-XL [58] adds ${\bm{y}}_{0}$ with the first frame of noisy input, which reduces information about ${\bm{y}}_{0}$ largely and thus mitigates the leakage issue. However, Fig. 3(a) indicates that the generated videos do not accurately preserve ${\bm{y}}_{0}$. Dynamicrafter [54] and SVD [9] directly feed ${\bm{y}}_{0}$ into I2V-DMs, enhancing detail preservation but also facilitating leakage. Dynamicrafter selectively refines spatial layers while preserving the pre-trained temporal layers, which contain motion priors and thus maintain motion dynamics to a certain extent. However, it does not inherently solve the leakage issue. SVD [9] introduces an additional motion score to compel the model to generate videos with motion to align with motion score, thereby reducing reliance solely on ${\bm{y}}_{0}$ and mitigating conditional image leakage. However, we argue that it does not address the fundamental challenge facing I2V-DMs, where I2V-DMs should generate clean video primarily from noisy input. An example in Fig. 3(b) highlights this limitation. By increasing the motion score, the method produces videos with significant camera movement to meet the scalar motion score but leaves objects static. In contrast, the strategy we propose in Sec. 4.1 results in a video with vivid and naturalistic movement, where both the people and the teapot exhibit dynamic motion, offering a more effective solution to the fundamental challenges of I2V-DMs. Furthermore, if an inappropriate motion score is provided during inference, the resulting video may exhibit unnatural motion or poor temporal consistency to match the given score (see Fig. 12 in the Appendix). In summary, despite efforts by these works to somewhat mitigate conditional image leakage, the issue persists, underscoring the ongoing need for further exploration.
Figure 4: Benefits of tuning $M$ and Analytic-Init w.r.t. visual quality. An appropriate $M$ can enhance performance by increasing motion without compromising other performance. A too-small $M$ delivers poor visual quality due to the training-inference gap. Analytic-Init helps to mitigate it.
4 Solving Conditional Image Leakage in Image-to-video Diffusion Models
In this section, we tackle the issue of conditional image leakage from two perspectives: first, by outlining a training-free inference strategy to mitigate it, and second, by presenting a training strategy designed to rectify this phenomenon during the training phase.
4.1 Inference Strategy
@@
As discussed in Sec. 3.1, conditional image leakage commonly occurs at large time steps. To address it, a straightforward solution is to start the generation process from an earlier time step $M\in[0,T]$, avoiding the unreliable later stages of I2V-DMs. Let $p_{M}(X_{M})$ denote the initial noise distribution at the start time $M$. Initially, we set $p_{M}(X_{M})=p_{T}(X_{T})$, i.e. ${\mathcal{N}}({\bm{0}},{\bm{I}})$ in VP-SDE [23] or ${\mathcal{N}}({\bm{0}},\sigma_{T}^{2}{\bm{I}})$ in VE-SDE [28]. As illustrated in the third column ($M=0.92T$) of Fig. 4, this straightforward strategy markedly improves motion dynamics without sacrificing other performance. However, a smaller $M$ value (e.g., $M=0.8T$) hurts the visual quality due to the training-inference discrepancy.
@@
To bridge the training-inference gap, we propose to minimize the KL divergence between the initial noise distribution $p_{M}(X_{M})$ and the actual marginal distribution of the diffusion process at time $M$, denoted as $q_{M}(X_{M})$. Inspired by existing work [4, 3, 55], when $p_{M}(X_{M})$ is modeled as a normal distribution ${\mathcal{N}}(X_{M};{\bm{\mu}}_{p},\sigma_{p}^{2}{\bm{I}})$, we derive analytical expressions for the optimal mean ${\bm{\mu}}_{p}^{*}$ and variance $\sigma_{p}^{2\ast}$ as follows. We refer to it as the Analytic Noise Initialization (Analytic-Init).
@@
Appendix A provides detailed proof of Proposition 1. Empirically, we can estimate ${\bm{\mu}}_{p}^{*}$ and $\sigma_{p}^{2\ast}$ using the method of moments [4]. Compared with the naive approach where $p_{M}(X_{M})=p_{T}(X_{T})$, Analytic-Init can achieve superior visual quality by reducing the training-inference gaps, especially with a smaller $M$. The theoretical insights and Proposition 1 are validated by the results in Tab. 4, measured by the commonly used FVD and IS metrics.
Figure 5: Effects of tuning $\mu(t)$ and $\beta_{m}$ in TimeNoise. Adjusting these parameters allows for a flexible trade-off between dynamic motion and image alignment. Higher noise levels result in more pronounced motion but reduced image alignment.
Figure 6: Effectiveness of TimeNoise to address conditional image leakage. We replicate the experiments described in Sec. 3.1 to evaluate TimeNoise. Applying TimeNoise ensures that $\hat{X}_{t\rightarrow 0}$ maintains motion dynamics comparable to the GT across all time steps, indicating the absence of conditional image leakage. Consequently, we can generate videos with high motion starting from $T$.
4.2 Training Strategy
The previous section explores strategies to mitigate conditional image leakage during inference. This section presents methods to address this issue during the training phase. As noted in Sec. 3.1, the conditional image ${\bm{y}}_{0}$ remains informative of the target video, leading I2V-DMs to depend excessively on it. Therefore, an intuitive idea is to perturb ${\bm{y}}_{0}$ to relieve the reliance on it.
Our first attempt is to introduce noise similar in scale to that used in $X_{t}$. This aims to equalize the model’s challenge in predicting clean video from $X_{t}$ or ${\bm{y}}_{t}$, reducing its dependency on ${\bm{y}}_{0}$. However, this strategy also makes it difficult to harness ${\bm{y}}_{0}$, resulting in lower video quality. To overcome this, we propose a noise distribution on ${\bm{y}}_{0}$ that introduces substantial noise to prevent leakage while maintaining a cleaner ${\bm{y}}_{0}$ to aid content generation. As discussed earlier, as time progresses, $X_{t}$ becomes less informative about $X_{0}$, increasing the risk of leakage. This insight leads us to develop a time-dependent noise distribution $p_{t}(\beta_{s})$ (TimeNoise). The design principle is to favor high noise levels at large time steps to sufficiently disrupt ${\bm{y}}_{0}$, shifting towards lower noise levels as the time step decreases. To achieve this, we employ a logit-normal distribution [18, 1] defined as below:

### v2

@@ Abstract
Diffusion models have obtained substantial progress in image-to-video generation. However, in this paper, we find that these models tend to generate videos with less motion than expected. We attribute this to the issue called conditional image leakage, where the image-to-video diffusion models (I2V-DMs) tend to over-rely on the conditional image at large time steps. We further address this challenge from both inference and training aspects. First, we propose to start the generation process from an earlier time step to avoid the unreliable large-time steps of I2V-DMs, as well as an initial noise distribution with optimal analytic expressions (Analytic-Init) by minimizing the KL divergence between it and the actual marginal distribution to bridge the training-inference gap. Second, we design a time-dependent noise distribution (TimeNoise) for the conditional image during training, applying higher noise levels at larger time steps to disrupt it and reduce the model’s dependency on it. We validate these general strategies on various I2V-DMs on our collected open-domain image benchmark and the UCF101 dataset. Extensive results show that our methods outperform baselines by producing higher motion scores with lower errors while maintaining image alignment and temporal consistency, thereby yielding superior overall performance and enabling more accurate motion control. The project page: https://cond-image-leak.github.io/.
@@ 1 Introduction
Image-to-video (I2V) generation aims to generate videos with dynamic and natural motion while maintaining the content of the given image. It allows users to guide video creation from the input image (and optional text), thus increasing controllability and flexibility in content creation. Like the remarkable progress in text-to-image (T2I) generation  [42, 39, 21, 8, 18, 5] and text-to-video (T2V) generation  [10, 46, 6, 11], diffusion models have also obtained promising results for I2V generation [9, 15, 63, 12, 68, 69]. However, such models are not fully understood.
Figure 1: The issue of existing I2V-DMs. Regardless of input motion scores (Input MS), the output motion scores (Output MS) are consistently lower than expected. In contrast, our method yields output motion scores either higher or lower than Input MS with reduced error.
@@
In this paper, we observe that existing image-to-video diffusion models (I2V-DMs) tend to generate videos with less motion than expected (see Fig. 1). We attribute this to a previously overlooked issue called conditional image leakage (see Sec. 3.1). Normally, the noisy input contains the motion information of the target video, and I2V-DMs should rely on it to predict motion, while the static conditional image provides content guidance. However, in practice, as the diffusion process progresses—especially at large time steps, the noisy input becomes heavily corrupted, while the conditional image preserves extensive detail of the target video. This biases the model to over-rely on the conditional image and neglect the noisy input, leading to videos with reduced motion. To validate this, we corrupt the ground truth (GT) clean video and compute one-step clean video predictions at each time step. As shown in Fig. 2, the predicted clean videos exhibit markedly reduced motion than GT at large time steps, indicating leakage.
@@
Based on the above analysis, we attempt to address the challenge from both inference and training aspects (see Sec. 3.2). First, we present a simple yet effective inference strategy that starts the video generation process from an earlier time step, thus avoiding the unreliable large-time steps of the I2V-DMs. To further enhance performance, we derive an initial noise distribution with optimal analytic expressions (Analytic-Init) by minimizing the KL divergence between it and the true marginal distribution to bridge the training-inference gap. Second, to mitigate leakage during training, we propose a time-dependent noise distribution (TimeNoise) that increases noise levels at larger time steps, effectively disrupting the conditional image and reducing model dependency on it. This is achieved by employing a logit-normal distribution with a center that gradually shifts over time. Finally, our method achieves higher motion scores with reduced motion score error and ensures that the predicted clean video maintains motion dynamics comparable to the ground truth across all time steps, effectively mitigating conditional image leakage. Notably, our general strategies are adaptable to various I2V-DMs based on both VP-SDE [12, 63, 69] and VE-SDE [9] framework.
@@
Empirically, we validate our methods on various I2V-DMs [9, 63, 12, 69] using our collected open-domain images (ImageBench) and the UCF101 dataset. We conduct a user study on ImageBench and report FVD [55], IS [7], and motion score [9] on UCF101. For motion-conditioned models [9, 69], we also report the motion score error between the generated video and the input motion score at different levels. Extensive experimental results demonstrate that our strategies outperform baselines by producing higher motion scores with lower errors while maintaining image alignment and temporal consistency, thereby yielding superior overall performance and enabling more accurate motion control.
@@ 2 Background
Diffusion models gradually perturb the data ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$ via a forward diffusion process and reverse the process to recover it. The forward transitional kernel $q_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})$ is given by
@@
where $\alpha_{t}$ and $\sigma_{t}$ are the noise schedule chosen to ensure that ${\bm{x}}_{T}$ contains minimal information about ${\bm{x}}_{0}$. Such forward diffusion processes can be viewed as stochastic differential equations (SDEs), among which two prevalent types are commonly used [48, 29]. One is the variance-preserving SDE (VP-SDE) [42, 24], where $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ with $\alpha_{t}\rightarrow 0$ as $t\rightarrow T$, ensuring $p_{T}({\bm{x}}_{T})=\mathcal{N}({\bm{x}}_{T};{\bm{0}},{\bm{I}})$. The other is the variance exploding SDE (VE-SDE), where $\alpha_{t}\equiv 1$ and $\sigma_{T}$ is set to a large constant, resulting in $p_{T}({\bm{x}}_{T})\approx{\mathcal{N}}({\bm{x}}_{T};{\bm{0}},\sigma_{T}^{2}{\bm{I}})$. Such models can be parameterized with a noise-prediction model $\epsilon_{\theta}({\bm{x}}_{t},t)$ (${\bm{\epsilon}}$-prediction) [24],and the parameters are learned by minimizing:
@@
where the noisy input ${\bm{x}}_{t}\sim q_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})$. Alternative parametrizations such as ${\bm{x}}_{0}$-prediction [29], ${\bm{v}}$-prediction [45], $F$-prediction [28] are also commonly applied. The ${\bm{\epsilon}}$-prediction and ${\bm{x}}_{0}$-prediction aims to predict the added noise or clean video from noisy input ${\bm{x}}_{t}$. Starting from ${\bm{x}}_{T}\sim p_{T}({\bm{x}}_{T})$, various samplers [47, 33, 34, 28] can be employed to generate data.
@@
Given an image ${\bm{y}}_{0}$ from the open domain, the goal of I2V is to generate a video $X_{0}=\{{\bm{x}}_{0}^{i}\}_{i=1}^{N}$ with dynamic and natural motion while keeping alignment with the appearance of ${\bm{y}}_{0}$. This task can be formulated as designing a conditional distribution $p_{\theta}(X_{0}|{\bm{y}}_{0})$, which is achieved by a conditional diffusion model minimizing:
@@
where $X_{t}\sim q_{t|0}(X_{t}|X_{0})$. Typically, ${\bm{y}}_{0}$ is the first frame of $X_{0}$ and DynamiCrafter [63] adopts a randomly selected frame from $X_{0}$ as ${\bm{y}}_{0}$. The key issue is to effectively integrate the conditional image ${\bm{y}}_{0}$ into the diffusion model. Most methods use CLIP image embeddings [40] to maintain the semantic content of ${\bm{y}}_{0}$. Notably, VideoCrafter1 [12] and Dynamicrafter [63] employ the last layer’s full patch visual tokens from the CLIP ViT, enriching the encoded information, and other approaches prefer the class token layer. Yet, solely depending on these embeddings, such as in VideoCrafter1 [12], compromises detail retention, resulting in degradation of the image alignment. To enhance detail representation, I2VGen-XL [68] combines the conditional image with the noisy initial frame, while VideoComposer [57] develops a STC-encoder for multiple conditions. Although superior to CLIP image embeddings, these strategies still fail to fully retain the conditional image content. To mitigate this, AnimateAnything [15], Dynamicrafter [63] and SVD [9] directly concatenate noisy video $X_{t}$ with ${\bm{y}}_{0}$, which injects detailed information to the model. Apart from the prior work mentioned before, we discuss other related work about diffusion models for image generation and video generation in the Appendix F.
Figure 2: Identifying conditional image leakage. As time step progresses, the noisy input becomes heavily corrupted, whereas the conditional image retains considerable detail from GT. This biases the model to over-rely on the conditional image at large $t$, resulting in videos with less motion than GT.
3 Method
Although existing I2V-DMs discussed in Sec. 2 have achieved significant progress, such models are not fully understood. In this section, we first identify a critical yet previously overlooked issue in I2V-DMs: conditional image leakage (CIL) (see Sec. 3.1). We then address this issue from both inference and training aspects accordingly (see Sec. 3.2). Finally, we offer insights into existing I2V-DMs through the lens of CIL (see Sec. 3.3).
@@ 3.1 Identifying Conditional Image Leakage in Image-to-video Diffusion Models
As shown in Fig. 1, we observe that, regardless of the input motion scores, the motion scores of generated videos from existing I2V-DMs [9, 69] are consistently lower than expected. This raises the question: why do these models always produce lower motion scores, rather than fluctuating above or below the expected values, as observed in our method?
To understand this, we need to consider the source of motion information in the generated videos, which comes from the noisy input $X_{t}$. Ideally, I2V-DMs should rely primarily on $X_{t}$ for motion, with the static conditional image ${\bm{y}}_{0}$ providing content guidance. However, as shown in Fig. 2, at large time steps, the noisy input $X_{t}$ becomes increasingly corrupted, while the conditional image ${\bm{y}}_{0}$ retains significant information of the target video. This biases the model to over-rely on the conditional image and neglect the noisy inputs, leading to videos with reduced motion.
To validate this, we corrupt a ground truth (GT) clean video $X_{0}$ via the forward transition kernel in Eq. (1) and use it as the noisy input to compute the one-step prediction $\hat{X}_{t\rightarrow 0}$ at time $t$:
@@
Ideally, $\hat{X}_{t\rightarrow 0}$ should predict the GT $X_{0}$ from noisy input $X_{t}$ and exhibit comparable motion dynamics. However, as shown in Fig. 2, as time progresses—particularly at large time steps, $\hat{X}_{t\rightarrow 0}$ exhibits markedly reduced motion than GT, indicating the conditional image leakage. This results in videos with reduced motion starting from time $T$. Notably, recent techniques that adjust the noise schedule towards higher noise levels [14, 27, 9, 32] may further exacerbate this issue (see Appendix C for details).
Figure 3: Benefits of Analytic-Init. (a) An early start time $M$ enhances motion but a too-small $M$ degrades visual quality due to the training-inference gap, which Analytic-Init helps to reduce. (b) Analytic-Init produces higher motion scores with lower errors, mitigating conditional image leakage.
3.2 Solving Conditional Image Leakage in Image-to-video Diffusion Models
Building upon the above analysis, this section presents general strategies to address the issue of conditional image leakage in both the inference and training aspects.
Inference strategy.
@@ 3.2 Solving Conditional Image Leakage in Image-to-video Diffusion Models
As discussed in Sec. 3.1, conditional image leakage easily occurs at large time steps. To this end, a straightforward solution is to start the generation process from an earlier time step $M\in(0,T)$, thus avoiding the unreliable later stages of I2V-DMs. Let $p_{M}(X_{M})$ denote the initial noise distribution at the start time $M$. Initially, we set $p_{M}(X_{M})=p_{T}(X_{T})$, i.e. ${\mathcal{N}}({\bm{0}},{\bm{I}})$ in VP-SDE [24, 63, 12, 69] or ${\mathcal{N}}({\bm{0}},\sigma_{T}^{2}{\bm{I}})$ in VE-SDE [29, 9]. As illustrated in Fig. 3 (a), this straightforward strategy markedly improves motion dynamics without sacrificing other performance. However, a smaller $M$ value (e.g., $M=0.8T$) results in poor visual quality due to the training-inference discrepancy.
@@
To mitigate this gap, we propose Analytic Noise Initialization (Analytic-Init) to refine the initial noise distribution $p_{M}(X_{M})$ by minimizing the KL divergence between it and the true marginal distribution $q_{M}(X_{M})$ of the forward diffusion process. Inspired by previous work [4, 3, 64], we demonstrate that when $p_{M}(X_{M})$ is modeled as a normal distribution ${\mathcal{N}}(X_{M};{\bm{\mu}}_{p},\sigma_{p}^{2}{\bm{I}})$, the optimal mean ${\bm{\mu}}_{p}^{*}$ and variance $\sigma_{p}^{2\ast}$ have analytical solutions, as stated in Proposition 1.
(a) TimeNoise $p_{t}(\beta_{s})$ with hyperparameters $\beta_{m}$ (maximum noise) and $\mu(t)$ (distribution center).
(b) Distribution center $\mu(t)=2t^{a}-1$ with $a$ controlling monotonic behavior flexibly.
(c) Effects of tuning $a$ in $\mu(t)$ and $\beta_{m}$, with higher noise levels enhancing motion but reducing alignment.
Figure 4: Visualization of TimeNoise and the impact of tuning its hyperparameters. (a) The designed $p_{t}(\beta_{s})$ favors high noise levels at large $t$, gradually shifting to lower noise levels as $t$ decreases. This is achieved by (b) $\mu(t)$ increasing monotonically with $t$. Finally, (c) modifying $a$ and $\beta_{m}$ enables a trade-off between dynamic motion and image alignment.
@@
The proof of Proposition 1 is provided in Appendix A. Empirically, ${\bm{\mu}}_{p}^{*}$ and $\sigma_{p}^{2\ast}$ can be estimated using the method of moments [4]. The steps for inference are outlined in Algorithm 1. As demonstrated by the qualitative results in Fig. 3 (a) and quantitative results in Tab. 8, Analytic-Init improves video quality by reducing the training-inference gap, especially for smaller $M$. Finally, as shown in Fig. 3 (b) and Tab. 1, Analytic-Init produces higher motion scores with lower errors, allowing for more accurate motion control and reducing conditional image leakage.
Figure 5: Benefits of TimeNoise. TimeNoise effectively reduces conditional image leakage by (a) generating $\hat{X}_{t\rightarrow 0}$ that maintains motion dynamics comparable to the GT across all time steps, and (b) achieving higher motion scores with lower errors.
Training strategy.
In this section, we show how to address the issue of conditional image leakage during the training phase. As outlined in Sec. 3.1, the conditional image ${\bm{y}}_{0}$ retains substantial details of the target video, causing I2V-DMs to rely heavily on it. To mitigate this, a natural approach is to perturb ${\bm{y}}_{0}$ to relieve this dependency. Our first attempt is to introduce noise at a similar scale to that in $X_{t}$, aiming to balance the model’s challenge of predicting clean video from $X_{t}$ or ${\bm{y}}_{0}$, thereby lessening reliance on ${\bm{y}}_{0}$. However, this strategy also makes it difficult to employ ${\bm{y}}_{0}$, resulting in lower video quality.
To overcome this, we propose a noise distribution on ${\bm{y}}_{0}$ that introduces substantial noise to prevent leakage while maintaining a cleaner ${\bm{y}}_{0}$ to aid content generation. Given that $X_{t}$ contains less information about $X_{0}$ as time progresses, increasing the risk of leakage, we further develop a time-dependent noise distribution $p_{t}(\beta_{s})$ (TimeNoise). The key principle is to favor high noise levels at large time steps to sufficiently disrupt ${\bm{y}}_{0}$, shifting towards lower noise levels as the time step decreases. To achieve this, we employ a logit-normal distribution [18, 1] defined as below: